find the equation of the tangent line to the graph of f at the point. write the eqUation in stabdard form.
4]od PxNx5, (242 94exy = Ge-354 1(0x4 2), 6-3) ६४: । (-1, 26) ६८४] =V+2x18, (14 NaxyTT (५-५४+४), (42)
Answer #1

2) febrer (2) = (-1) Equation of Tangent line : y Y = m (-) where m= f(n) at (211) (61) m = f (2) +840 674) (134) = -32 - m=